# Bound of Casimir Effect by Holography

### v2

@@ Abstract
Higher dimensions — Let us comment on the holographic bound of Casimir effect for a strip in higher dimensions. The corresponding Casimir amplitude $\kappa_{1}$ and norm of displacement operator $C_{D}$ is given by Miao:2024ddp for Einstein gravity. By taking the limit $T\to-(d-1)$, we derive a universal lower bound
$(\frac{-\kappa_{1}}{C_{D}})\geq\frac{-2^{d-2}d\pi^{d-\frac{1}{2}}\Gamma\left(\frac{d-1}{2}\right)\Gamma\left(\frac{1}{d}\right)^{d}}{\Gamma(d+2)\left(d\Gamma\left(\frac{1}{2}+\frac{1}{d}\right)\right)^{d}}.$ (17)
[…]
Non-CFTs — Interestingly, the lower bound of (3) also works for a general class of non-CFTs. Take an example of a 3d free scalar with mass $m$ and coupling $\xi$ with the Ricci scalar. It is a CFT only if $\xi=1/8$ and $m=0$. For non-CFT parameters, the energy density between parallel plates is no longer a constant Bordag:2009zz . What is worse, it is divergent near the boundary. Fortunately, the total energy per area is finite Bordag:2009zz . On the other hand, the pressure is still a finite constant. Thus, we define $\kappa_{1}=\langle T_{nn}(0)\rangle L^{3}/(-2)$ generally. By applying the textbook method, we obtain for both RBC and DBC, i.e., $16\pi\kappa_{1}=-2M^{2}\log\left(1-e^{-2M}\right)+2M\text{Li}_{2}\left(e^{-2M}\right)+\text{Li}_{3}\left(e^{-2M}\right)$, where $M=mL>0$ is a dimensionless parameter, and Li denotes the Polylogarithmic function. Remarkably, $\kappa_{1}$ is independent of $\xi$ and decreases with mass
@@
It means the Casimir effect is suppressed by mass Bordag:2009zz . For non-CFTs, we can define $C_{D}$ via $\langle D(y)D(0)\rangle=C_{D}/|y|^{6}+O(1/|y|^{5})$ in half space, where $D=T_{nn}(0)$ Billo:2016cpy . At small distance $m|y|\ll 1$, the mass effect can be ignored. As a result, $C_{D}$ is independent of mass. The above discussions show $(-\kappa_{1}/C_{D})$ is larger than that of massless scalar, thus obeys the holographic lower bound (3). The above arguments apply to general massive free field theories.
Discussions — In conclusion, we propose that holography sets a fundamental bound for the Casimir amplitude ratio to the norm of displacement operator. The Casimir amplitude is generally defined as the average pressure $\bar{t}_{nn}$ on the boundary. We take 3d CFTs in a strip as an example and then briefly discuss the generalization to other boundary shapes, higher dimensions, and non-CFTs. It is interesting to test our proposal in experiments of critical systems Krech:Casimir Effect . Finding a field-theoretical proof or counterexample for the holographic bound of the Casimir effect is also enjoyable. Usually, a bound is set by either free or strongly coupled theory. For entanglement entropy (EE), there is a similar bound set by free theory, i.e., $(-F_{0}/C_{T})\geq(-F_{0}/C_{T})_{\text{free scalar}}$ Bueno:2023gey , where $F_{0}$ is the universal coefficient of EE and $C_{T}$ is a bulk central charge. We remark that the dual brane for EE is tensionless while the one for the Casimir effect is tensive. With a tensive brane, we have more parameters to get a smaller lower bound of Casimir effect in holography than in free theory. Ising model of Table. 1 also shows a free scalar does not set the lower bound of Casimir effect. In AdS3, conformal bootstrap restricts brane tension $|T|\leq 0.99$ sightly stronger than $|T|\leq 1$ Collier:2021ngi . If this is also the case for both $T$ and DGP coupling $\lambda$ in higher dimensions, it may give a slightly stronger lower bound of the Casimir effect. Anyway, this letter takes a significant step toward exploring the lower bound of the Casimir effect. Many open problems remain to be explored.
@@ Acknowledgements
We thank T. Takayanagi, J. X. Lu for valuable comments and discussions. We are grateful to the conference “Gauge Gravity Duality 2024”, where the work was completed. This work is supported by the National Natural Science Foundation of China (No.12275366).
@@ Appendix A Bound of Casimir effect from DGP gravity
This appendix studies the holographic bound of Casimir effect in DGP gravity. In particular, we give detailed derivations of eqs.(8, 10) in the main text.
Let us quickly recall some key points. The vacuum of a strip is dual to AdS soliton
@@
where $h(z)=1-z^{3}/z_{h}^{3}$, and $\theta$ has the range $0\leq\theta\leq L$ on the AdS boundary $z=0$. We label the embedding function of brane $Q$ as
@@
where the tension is parameterized as Miao:2023mui
@@
This parameterization can be derived from the gravity dual of half space. Since the region near any smooth boundary can be approximately by half space at leading order of $z$, it works for general boundaries. Substituting (20) into NBC (21), we derive near the two boundaries of strip
@@
Let us first study the case with negative brane tension $T\leq 0$, the other case with $T>0$ can be obtained by the complement of the first case. This is the typical trick used in Fujita:2011fp . Let us explain more. See Fig. 2 for the geometry of holographic dual of strip. The region between red/blue curves (branes) and black line is the bulk dual of strip I with $T\leq 0$; its complement in bulk is the gravity dual for strip II (green line) with $T>0$. As shown in Fig. 2, the gravity duals of strip I and strip II shares the same EOW branes (blue or red curve depending on the theory parameter). As a result, the NBCs (21) for strip I and strip II take the same values. However, the extrinsic curvatures $K_{ij}$ flip signs, while the induced metric $h_{ij}$ and intrinsic Ricci tensor $\mathcal{R}_{ij}$ remain invariant when crossing the branes. In the viewpoint of NBC (21), $T$ and $\lambda$ equivalently change signs when crossing the branes. Note that the bulk dual for strip II contains the ‘horizon’ $z=z_{h}$. To remove the conical singularity at $z=z_{h}$, we fix the period of angle $\theta$ as $\beta=4\pi/|h^{\prime}(z_{h})|=4\pi z_{h}/3$. Thus for $T>0$, the left and right vertical dotted lines of Fig. 2 are identified due to the periodicity of angle $\theta$. On the other hand, for strip I with $T\leq 0$, the ‘horizon’ and potential conical singularity is hidden behind the branes (red/blue curves). Thus, the conical singularity is irrelevant and $\theta$ can be non-periodic for $T\leq 0$.
Without loss of generality, we focus on the strip I with $T\leq 0$ and $z_{h}=1$ below. According to (23), the blue and red curves of Fig. 2 correspond to $\rho<0$ and $\rho>0$, respectively. We focus on the left half of branes, the right half can be obtained by symmetry. NBC (21) yields an independent equation for $T\leq 0$
@@
Let us verify that it corresponds to the left half of branes. From (22,24), we derive $S^{\prime}(0)=-\sinh(\rho)$ near $\theta_{1}$ of Fig. 2, which agrees with the first equation of (23). As shown in Fig. 2, $S^{\prime}(z_{\text{max}})=\infty$ and $S^{\prime}(z_{c})=0$ at the turning points $z_{\text{max}}$ and $z_{c}$ for the left half of branes. Substituting $S^{\prime}(z_{\text{max}})=\infty$ and $S^{\prime}(z_{c})=0$ into (24), we obtain
@@
where $H=\frac{1}{2}h(z)\left(1+2\lambda\sqrt{h\left(z_{\max}\right)}+X\right)$, $X=\sqrt{1+4\lambda^{2}h(z)-4\lambda\sqrt{h\left(z_{\max}\right)}}$. We choose the positive sign, i.e., $S^{\prime}(z)\sim\Big{(}\sqrt{h\left(z_{\max}\right)}-\lambda h(z)\Big{)}$, for the left halves of red and blue curves to get the correct behavior $S^{\prime}(0)=-\sinh(\rho)$ near $\theta_{1}$ of Fig. 2. The negative sign of (26) corresponds to the right halves of branes. From (20,26), we derive the width of strip I for $T\leq 0$
@@
where $z_{\max}=(1-T^{2}/4)^{1/3}$ from (25). We stress that the above formula works for both the cases of blue curve and red curve of Fig. 2. In particular, the integrand $2S^{\prime}(z)\sim 2\Big{(}\sqrt{h\left(z_{\max}\right)}-\lambda h(z)\Big{)}$ flips signs at turning point $z_{c}$ (25), which is the expected feature for the red curve. It suggests the integrand cannnot be chosen as the absolute value $2|S^{\prime}(z)|\sim 2\Big{|}\sqrt{h\left(z_{\max}\right)}-\lambda h(z)\Big{|}$, which would miss the case of red curve.
@@
One can check $L$ is continuous at $T=0$, which is a test of our results. See Fig. 3 for example. As mentioned in main text, the constraint $0\leq L$ set an upper bound of $\lambda$. It is clear from (27) that large $\lambda$ would give negative $L$. We numerically get the range $0\leq\lambda\leq 1$ for $L\geq 0$. As shown in Fig. 3 , the case $20\leq\lambda\leq 1/$ 2 can continuously transform into that of Einstein gravity, while the case $1/2<\lambda\leq 1$ cannot. We name them normal and singular phases, respectively. $\lambda=1/2$ is the phase-transition point.
Figure 3: Strip widths in normal phase with $20\leq\lambda<1/$ 2 (up) and in singular phase with $1/2<\lambda\leq 1$ (down). Note that $L$ can be larger than $\beta$ for negative enough $\rho$ in the normal phase. We show only range of $L\leq\beta$ for simplicity. On the other hand, $L$ is always smaller than $\beta$ in the singular phase. In both cases, $0\leq L$ imposes an upper bound of $\rho$.
@@
[…]
@@
The norm of displacement operator becomes for $\rho<0$
@@
Interestingly, the second term of (A) vanishes for $\lambda\to 1/2$. In fact, this is also the case for higher-order expansions. From (33) and $L_{\text{I}}=\beta=4\pi/3$, we solve

### v3

@@ Abstract
$(\frac{-\kappa_{1}}{C_{D}})|_{d=2}\geq-\frac{\pi^{3}}{12}.$ (19)
According to Bloete:1986qm , the inequality is saturated for general 2D CFTs with the same boundary conditions on the two planes. On the other hand, if we impose different boundary conditions on the two planes, we get larger Casimir energy Bachas:2006ti ; Diatlyk:2024qpr
$(-\kappa_{1})_{\text{differ BC}}\geq(-\kappa_{1})_{\text{same BC}}.$ (20)
Note that $C_{D}=c/(2\pi^{2})$ is independent of the choices of boundary conditions, where $c$ is the central charge of 2d CFTs. Considering all the above statements, we get solid proof of the bound (19) for 2D CFTs, which is a strong support for our proposal (18).
Non-CFTs — Interestingly, the lower bound of (3) also works for a general class of non-CFTs. Take an example of a 3D free scalar with mass $m$ and coupling $\xi$ with the Ricci scalar. It is a CFT only if $\xi=1/8$ and $m=0$. For non-CFT parameters, the energy density between parallel plates is no longer a constant Bordag:2009zz . What is worse, it is divergent near the boundary. Fortunately, the total energy per area is finite Bordag:2009zz . On the other hand, the pressure is still a finite constant. Thus, we define $\kappa_{1}=\langle T_{nn}(0)\rangle L^{3}/(-2)$ generally. By applying the textbook method, we obtain for both RBC and DBC, i.e., $16\pi\kappa_{1}=-2M^{2}\log\left(1-e^{-2M}\right)+2M\text{Li}_{2}\left(e^{-2M}\right)+\text{Li}_{3}\left(e^{-2M}\right)$, where $M=mL>0$ is a dimensionless parameter, and Li denotes the polylogarithmic function. Remarkably, $\kappa_{1}$ is independent of $\xi$ and decreases with mass,
@@
For free non-CFTs, we can define $C_{D}$ via $\langle D(y)D(0)\rangle=C_{D}/|y|^{6}+O(1/|y|^{5})$ in half-space, where $D=T_{nn}(0)$ Billo:2016cpy . At a small distance $m|y|\ll 1$, the mass effect can be ignored. As a result, $C_{D}$ is independent of mass. The above discussions show $\kappa_{1}/C_{D}$ is larger than that of a massless scalar, thus obeys the holographic lower bound (3). The above arguments apply to general massive free theories. The Casimir effect arises from long-range quantum correlations Dantchev:2022hvy . When mass is introduced, these correlations become short range, which generally reduces the Casimir amplitude $\kappa_{1}$ Bordag:2009zz . Recall that $C_{D}$ is mass independent. Consequently, $\kappa_{1}/C_{D}$ reaches its minimum value in the massless limit for general free theories.
Discussions — In conclusion, we propose that holography sets a fundamental bound on the ratio of the Casimir amplitude to the displacement operator norm. The Casimir amplitude is defined as the average pressure $\bar{t}_{nn}$ on the boundary. We take 3D CFTs in a strip as an example and then briefly discuss the generalization to other boundary shapes, higher dimensions, and free non-CFTs. It is interesting to test our proposal in experiments of critical systems Krech:Casimir Effect . Finding a field-theoretical proof or counterexample for our proposal is also enjoyable. Our findings help estimate how much the Casimir effect contributes to dark energy and whether sufficient Casimir energy can be generated to support a traversable wormhole macroscopically. Additionally, it helps evaluate the maximum Casimir force between nanodevices, which can enhance nanotechnology designs. Usually, a bound is set by either strongly coupled or free theories Kovtun:2004de ; Hofman:2008ar ; Hofman:2016awc . For instance, the KSS bound Kovtun:2004de for fluid is set by the strongly coupled theory dual to gravity, while the Hofman-Maldacena bound Hofman:2008ar ; Hofman:2016awc for bulk central charges is set by free theories. For entanglement entropy, there is a similar bound set by free theories, i.e., $(-F_{0}/C_{T})\geq(-F_{0}/C_{T})_{\text{free scalar}}$ Bueno:2023gey , where $F_{0}$ is the universal coefficient of entanglement entropy and $C_{T}$ is a bulk central charge. We remark that the dual brane for entanglement entropy is tensionless while the one for the Casimir effect is tensive. With a tensive brane, we have more parameters to get a smaller lower bound of the Casimir effect in holography than in free theory. The Ising model of Table 1 also shows that a free scalar does not set the lower bound of the Casimir effect. In AdS3, a conformal bootstrap restricts brane tension $|T|\leq 0.99$ sightly stronger than $|T|\leq 1$ Collier:2021ngi . If this is also the case for both $T$ and DGP coupling $\lambda$ in higher dimensions, it may give a slightly stronger lower bound of the Casimir effect. Anyway, this Letter takes a significant step toward exploring the lower bound of the Casimir effect. Many open problems remain to be explored, particularly the generalization to non-CFTs with interactions.
@@ Acknowledgements
We thank T. Takayanagi and J. X. Lu for valuable comments and discussions. We are grateful to the conference “Gauge Gravity Duality 2024,”where the work was completed. This work is supported by the National Natural Science Foundation of China (No.12275366).
@@ Appendix A Bound of Casimir effect from DGP gravity
This supplemental material studies the holographic bound of the Casimir effect in DGP gravity. We give detailed derivations of equations (11, 12, 13) in the main text.
Let us quickly recall some key points. The vacuum of a strip is dual to the AdS soliton
@@
where $h(z)=1-z^{3}/z_{h}^{3}$, and $\theta$ has the range $0\leq\theta\leq L$ on the AdS boundary $z=0$. We label the embedding function of the brane $Q$ as
@@
where the tension is parameterized as
@@
This parameterization can be derived from the gravity dual of a half-space. Since the region near any smooth boundary can be approximated by a half-space at the leading order of $z$, it works for general boundaries. Substituting (23) into NBC (24), we derive near the two boundaries of strip
@@
Let us first study the case with negative brane tension $T\leq 0$; the other case with $T>0$ can be obtained by the complement of the first case. It is the typical trick used in Fujita:2011fp . Let us explain more. See Fig. 2 for the geometry of the holographic dual of the strip. The region between red/blue curves (branes) and the black line is the bulk dual of strip I with $T\leq 0$; its complement in the bulk is the gravity dual for strip II (green line) with $T>0$. As shown in Fig. 2, the gravity duals of strip I and strip II share the same EOW branes (blue or red curve depending on the theory parameter). As a result, the NBCs (24) for strip I and strip II take the same values. However, the extrinsic curvatures $K_{ij}$ flip signs, while the induced metric $h_{ij}$ and intrinsic Ricci tensor $\mathcal{R}_{ij}$ remain invariant when crossing the branes. From the viewpoint of NBC (24), $T$ and $\lambda$ equivalently change signs when crossing the branes. Note that the bulk dual for strip II contains the ‘horizon’ $z=z_{h}$. To remove the conical singularity at $z=z_{h}$, we fix the period of angle $\theta$ as $\beta=4\pi/|h^{\prime}(z_{h})|=4\pi z_{h}/3$. Thus, for $T>0$, the left and right vertical dotted lines of Fig. 2 are identified due to the periodicity of angle $\theta$. On the other hand, for strip I with $T\leq 0$, the ‘horizon’ and potential conical singularity is hidden behind the branes (red/blue curves). Thus, the conical singularity is irrelevant and $\theta$ can be non-periodic for $T\leq 0$.
Without loss of generality, we focus on the strip I with $T\leq 0$ and $z_{h}=1$ below. According to (26), the blue and red curves of Fig. 2 correspond to $\rho<0$ and $\rho>0$, respectively. We focus on the left half of the branes; the right half can be obtained by symmetry. NBC (24) yields an independent equation for $T\leq 0$
@@
Let us verify that it corresponds to the left half of the branes. From (25,27), we derive $S^{\prime}(0)=-\sinh(\rho)$ near $\theta_{1}$ of Fig. 2, which agrees with the first equation of (26). As shown in Fig. 2, $S^{\prime}(z_{\text{max}})=\infty$ and $S^{\prime}(z_{c})=0$ at the turning points $z_{\text{max}}$ and $z_{c}$ for the left half of branes. Substituting $S^{\prime}(z_{\text{max}})=\infty$ and $S^{\prime}(z_{c})=0$ into (27), we obtain
@@
where $H=\frac{1}{2}h(z)\left(1+2\lambda\sqrt{h\left(z_{\max}\right)}+X\right)$, $X=\sqrt{1+4\lambda^{2}h(z)-4\lambda\sqrt{h\left(z_{\max}\right)}}$. We choose the positive sign, i.e., $S^{\prime}(z)\sim\Big(\sqrt{h\left(z_{\max}\right)}-\lambda h(z)\Big)$, for the left halves of red and blue curves to get the correct behavior $S^{\prime}(0)=-\sinh(\rho)$ near $\theta_{1}$ of Fig. 2. The negative sign of (29) corresponds to the right halves of branes. From (23,29), we derive the width of strip I for $T\leq 0$
@@
where $z_{\max}=(1-T^{2}/4)^{1/3}$ from (28). We stress that the above formula works for both the cases of the blue curve and the red curve of Fig. 2. In particular, the integrand $2S^{\prime}(z)\sim 2\Big(\sqrt{h\left(z_{\max}\right)}-\lambda h(z)\Big)$ flips signs at turning point $z_{c}$ (28), which is the expected feature for the red curve. It suggests the integrand cannot be chosen as the absolute value $2|S^{\prime}(z)|\sim 2\Big|\sqrt{h\left(z_{\max}\right)}-\lambda h(z)\Big|$, which would miss the case of the red curve.
@@
One can check that $L$ is continuous at $T=0$, which is a test of our results. See Fig. 3 for example. As mentioned in the main text, the constraint $0\leq L$ sets an upper bound of $\lambda$. It is clear from (30) that a large $\lambda$ would give a negative $L$. We numerically get the range $0\leq\lambda\leq 1$ for $L\geq 0$. As shown in Fig. 3, the case $0\leq\lambda\leq 1/2$ can continuously transform into that of Einstein gravity, while the case $1/2<\lambda\leq 1$ cannot. We name them normal and singular phases, respectively. $\lambda=1/2$ is the phase-transition point.
Figure 3: Strip widths in normal phase with $0\leq\lambda<1/2$ (top) and in singular phase with $1/2<\lambda\leq 1$ (bottom). Note that $L$ can be larger than $\beta$ for negative enough $\rho$ in the normal phase. We show only range of $L\leq\beta$ for simplicity. On the other hand, $L$ is always smaller than $\beta$ in the singular phase. $0\leq L$ imposes a lower bound of $\rho$ in the singular phase, and a upper bound of $\rho$ in both phases.
@@
$L_{\text{I}}=\int_{0}^{1}dy\frac{2\sqrt[3]{1-2\lambda}}{\sqrt[6]{x}\sqrt{1-y^{3}}}+(1-2\lambda)\ O\left(\frac{x}{(1-2\lambda)^{2}}\right)^{\frac{5}{6}}$
@@
The norm of the displacement operator becomes for $\rho<0$
@@
Interestingly, the second term of (A) vanishes for $\lambda\to 1/2$. It is also the case for higher-order expansions. From (36) and $L_{\text{I}}=\beta=4\pi/3$, we solve